OpenStudy (anonymous):

Jas I am sorry, my browser had an error and I had to close it. I am still listening

OpenStudy (anonymous):

refresh ur page

OpenStudy (anonymous):

and do not type in ASK QUESTION COLOUMN type ur response under my response

OpenStudy (anonymous):

now that's right

OpenStudy (anonymous):

This is my first time on this website.

OpenStudy (anonymous):

oh it's ok

OpenStudy (anonymous):

The problem is multiply to complex fractions with a varible.

OpenStudy (anonymous):

I don't know how to express it using equation. It will not let me express it in fractions.

OpenStudy (anonymous):

k you write in words i will try to understand

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

two and two nineths * x, over four thirds * three halves

OpenStudy (anonymous):

4/3*2/3=8/9 (20*x/9)/8/9=((20*x)/9)*(9/8)=(20*x)/8=(5*x)/2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so any more q

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

two and two nineths * x, over four thirds * three halves

OpenStudy (anonymous):

express in numbers as i explained u

OpenStudy (anonymous):

I got for my answer as 2X over 18

OpenStudy (anonymous):

I wish I had a way of writing it better.

OpenStudy (anonymous):

you have 2/3 in your expression and it should have been 3/2

OpenStudy (anonymous):

The first side is 18/9 over 4/3

OpenStudy (anonymous):

This is then multiplied by X over 1 and 1/2. Is that clearer?

OpenStudy (anonymous):

sry i m busy with other prob so i will reply u after some time

OpenStudy (anonymous):

I am sorry. I am at work and had to step away.

OpenStudy (anonymous):

Maybe this would be better. 20 X _9_*___ 4 3 _____ 3 2

OpenStudy (anonymous):

i m not getting it

OpenStudy (anonymous):

(20/9 * X)/(4/3 * 3/2)

OpenStudy (anonymous):

there is two fractions on the top being multiplied and divided by two fractions on the bottom being multiplied.

OpenStudy (anonymous):

(20/9 * X) / (4/3 * 3/2)=

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

(20/9*x)/(2)=10/9*x